Use inductive reasoning to predict the most probable next number in the following lists.

a) 5, 10, 15, 20, 25, ?

b) 2, 5, 10, 17, 26, ?


1
Expert's answer
2022-05-11T09:21:53-0400

a. Study the sequence

We realize that we are supposed to add 5 to each number

5+5=10

10+5=15

15+5=20

20+5=25

25+5=30

Most probable next number =30


b. Study the sequence,

We realize that we are supposed to add odd numbers starting from 3 to the next.


2


2+3=5


5+5=10


10+7=17


17+9=26


26+11=37




Most probable next number =37